OpenStudy (anonymous):

one liter of paint covers 100 square feet. a wall that measures 32 feet by 16 feet is going to be painted. a. about how many gallons of paint will it take to paint the whole wall? b. if the paint can only be bought in gallons containers,how many containers will be purchased? c. how many liters of paint will be remaining in the container?

OpenStudy (anonymous):

32*16= 512 how much you need to paint

OpenStudy (campbell_st):

find the area of the wall in square feet Area = 32 x 16 next find how many litres of paint are needed No. of Litres = Area/100 convert the number of litres to gallons... you need to find the conversion 1 US gallon = 3.7854 litres 1 Imperial gallon = 4.546 liters For now many gallon cans, round your answer to the 1st part... number of gallons up to the nearest whole number. last part use the conversion again

OpenStudy (anonymous):

512/100 = 5.12 liters

OpenStudy (anonymous):

so 5.12 is for c?

OpenStudy (anonymous):

convert liters to gallons 1 litter= .264172 gallon